OpenStudy (anonymous):

point slope form from 4,4 to 6,8

OpenStudy (anonymous):

@ranga

OpenStudy (ranga):

Find the slope first. m = (y2-y1) / (x2-x1) Then point slope form is: (y-y1) = m(x-x1)

OpenStudy (anonymous):

8-4/6-4

OpenStudy (anonymous):

correct?

OpenStudy (ranga):

Use parenthesis. m = (8-4)/(6-4) = 4/2 = 2 Point slope form is: (y-y1) = m(x-x1) y - 4 = 2(x - 4)